OpenStudy (anonymous):

The flywheel of an engine has a moment of inertia 2.10{\rm kg} \cdot {\rm m}^{2} about its rotation axis. What constant torque is required to bring it up to an angular speed of 450rev/min in a time of 8.50s , starting from rest? What is its final kinetic energy?

OpenStudy (anonymous):

\[\Sigma \tau=I\alpha \\ \alpha=\frac{\Delta \omega}{\Delta t} \] Strategy: 1) Find alpha first 2) Find torque later :)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

how to find the Wo?

OpenStudy (anonymous):

It's zero (from rest)

OpenStudy (anonymous):

oh okay thanks ill try to solve now

OpenStudy (anonymous):

i got \[\alpha=47.123\]

OpenStudy (anonymous):

next is?

OpenStudy (anonymous):

Plug in this alpha to \[\Sigma \tau=I\alpha \] I=2.10

OpenStudy (anonymous):

i did it. but when i entered the answer in mastering physics its worng -_- N.m

OpenStudy (anonymous):

Omega must be in radians per second

OpenStudy (anonymous):

oh now i know whats wrong. i didnt divide it with time. waiiit :) ill see

OpenStudy (anonymous):

now correct :D how about the second question? how to do it :)

OpenStudy (anonymous):

\[K=\frac12 I\omega^2\]

OpenStudy (anonymous):

same value from the first? the w?

OpenStudy (anonymous):

Yes, final omega in rad/s

OpenStudy (anonymous):

okay 2x

OpenStudy (anonymous):

got new question. ill post now :)

OpenStudy (anonymous):

Keep posting :) Lot of people will help you here.